OpenStudy (anonymous):

A restaurant has one type of lemonade that is 12% sugar and another that is 7% sugar. How many gallons of each type does that restaurant need to make 20 gallons of a lemonade mixture that is 10% sugar? 12 gallons of the 12% lemonade and 8 gallons of the 7% lemonade. 10 gallons of the 12% lemonade and 10 gallons of the 7% lemonade. 8 gallons of the 12% lemonade and 12 gallons of the 7% lemonade. 2 gallons of the 12% lemonade and 18 gallons of the 7% lemonade.

OpenStudy (anonymous):

x = gallons of 12% sugar lemonade y = gallons of 7% sugar lemonade We can write 2 equations, now. x + y = 20.......That's the easier one. .12x + .07y = 0.1(20).....That's the other one 2 equations, 2 unknowns

OpenStudy (anonymous):

I'm not sure I follow. I mean I get where you got the equation. But they all equal 20.

OpenStudy (anonymous):

hello do u need help

OpenStudy (anonymous):

Yes, I do.

OpenStudy (anonymous):

well wat do u need help with

OpenStudy (anonymous):

I don't quite understand the question I posted above. It seems as if I need to combine it to get the answer but I'm not fully sure.

OpenStudy (anonymous):

that is egg zackuly wat u need to do

OpenStudy (anonymous):

i be right bac hav to feed my 2 dogs it will take about 4 to 5 min...

OpenStudy (anonymous):

They don't both equal 20. The 2nd one equals 0.1(20). That means 0.1 times 20...or 2. The 2nd equation equals 2.

OpenStudy (mathstudent55):

The way to solve a problem like this is by writing the two equations you were shown above. Then you need to solve the equations. Here's a little explanation that may be of some help. The first step is to pick two variables to represent the unknown amounts of the two lemonades. Let the amount of 12% lemonade be x. Let the amount of the 7% lemonade be y. Since 20 liters of lemonade are being made, x liters of 12% + y liters of 7% will equal 20 liters of 10%. That gives you the equation x + y = 20 Ok so far?

OpenStudy (anonymous):

Yeah, I'm understanding so far.

OpenStudy (anonymous):

im bac so do the variables first x+y -20

OpenStudy (anonymous):

x+y = 20

OpenStudy (mathstudent55):

Ok, since you have two unknowns, you need two equations. Now we turn to the amount of sugar in the three types of lemonade to find a second equation.

OpenStudy (mathstudent55):

A lemonade with 12% sugar has 0.12 times its amount in sugar. The x gallons of 12% lemonade contain 0.12x sugar. The y gallons of 7% lemonade contain 0.07y sugar. Ok?

OpenStudy (anonymous):

So it would be 0.12x + 0.07y = 20 right?

OpenStudy (mathstudent55):

No, but close. The amount of sugar is 0.12x and 0.07y from both lemonades. They are added together, so 0.12x + 0.07y is correct. The 10% resultant lemonade has a volume of 20 gal and a sugar concentration of 10%, so it has 0.10(20) sugar. So now you have your second equation: 0.12x + 0.07y = 0.10(20) or 0.12x + 0.07y = 2

OpenStudy (anonymous):

Ohh I see

OpenStudy (anonymous):

good but not quite, but you you have better help so im going to let him help you so good bye,talk to you latter if you still have questions just send me a message

OpenStudy (anonymous):

@DIANTE Oh well thanks, and I will c:

OpenStudy (anonymous):

k c ya

OpenStudy (mathstudent55):

Now you have your two equations: x + y = 20 0.12x + 0.07y = 2 Now you need to solve them simultaneously (as a system of equations).

OpenStudy (anonymous):

wait i have 1 question wat grade r u n

OpenStudy (anonymous):

im in the 9th

OpenStudy (anonymous):

@DIANTE I am in 9th also. @mathstudent55 Don't I have to use subtraction on both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its a 2 step equation

OpenStudy (mathstudent55):

Start by multiplying the entire second equation by 100 to get rid of the decimals.

OpenStudy (mathstudent55):

You get this: x + y = 20 12x + 7y = 200

OpenStudy (mathstudent55):

Now multiply the first equation by -7. Then add it to the second equation.

OpenStudy (anonymous):

mathstudent55 u r good at this wat grade r u in

OpenStudy (mathstudent55):

-7x - 7y = -140 (first eq multiplied by -7) 12x + 7y = 200 -------------- (now add equations) 5x = 60 Solve for x: x = 12 Now use the first original equation to solve for y: x + y = 20 12 + y = 20 y = 8 The answer is 12 gal 12% lemonade + 8 gal 7% lemonade.

OpenStudy (anonymous):

Oh now I understand how you got it. Thank you so much!

OpenStudy (anonymous):

what school do u go to

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Connections

OpenStudy (anonymous):

in GA

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

I live in GA I go 2 Langston Hughes High school

OpenStudy (anonymous):

Oh that's cool

OpenStudy (anonymous):

i no right

OpenStudy (anonymous):

well im bout to help someone else

OpenStudy (anonymous):

That Connexus Cheating.